# Unifying Correspondence, Pose and NeRF for Pose-Free Novel View Synthesis from Stereo Pairs

### 12 Dec 2023 v1

@@ 2 Related Work
3 Unified Framework for Pose-Free Novel View Synthesis
@@ 3.1 Problem Formulation
Assuming an unposed pair of images $I_{1},I_{2}\in\mathbb{R}^{H\times W\times 3}$ taken from different viewpoints as the input, our goal is to synthesize an image $\hat{I}_{t}$ from a novel view. Different from classical generalized NeRF tasks [51, 9, 59], our task is additionally challenged by the absence of camera pose between the input images. To this end, we estimate the relative camera pose between $I_{1},I_{2}$ as ${P}_{2\leftarrow 1}\in\mathbb{R}^{4\times 4},$ consisting of rotation $R\in\mathbb{R}^{3\times 3}$ and translation $T\in\mathbb{R}^{3\times 1}$, and deduce $P_{2\leftarrow t}=P_{2\leftarrow 1}P_{1\leftarrow t}$ with $P_{1\leftarrow t}$ as the desired rendering viewpoint, which are then used in conjunction with the extracted feature maps to compute the pixel color at the novel view by the renderer.
@@ 3.2 Cost Volume Construction
Unlike the previous methods that only consider local receptive fields within their cost volumes [56, 9, 27, 11], we consider all pairs of similarities between features to handle both small and large displacements. Specifically, we compute and store the pairwise cosine similarity between features, obtaining a 4D cost volume $\{C^{l}\}_{l=1}^{L}\in\mathbb{R}^{h^{l}\times w^{l}\times h^{l}\times w^{l}}$, where $L$ is the number of levels.
@@ 3.4 Flow and Relative Pose Estimation
Making use of the cost volume from previous steps, we define a dense flow field, $F(i)$ that warps all pixels in image $I_{1}$ towards $I_{2}$. We also estimate relative camera pose $P_{2\leftarrow 1}$ from this cost volume as it sufficiently embodies confidence scores and spatial correspondences [8, 40]. To estimate the dense flow map $F_{2\leftarrow 1}$, we can simply apply argmax to find the highest scoring correspondences. While this may be sufficient for image pairs with large overlapping regions, we account for the potential occlusions by computing a confidence map. Specifically, following [34], we obtain a cyclic consistency map $M_{2\leftarrow 1}(i)$ using the reverse field $F_{1\leftarrow 2}$ as an additional input, and check if the following condition is met for consistency: $||F_{2\leftarrow 1}(i)+F_{1\leftarrow 2}(i+F_{2\leftarrow 1}(i))||_{2}<\tau$, where $||\cdot||_{2}$ denotes frobenius norm and $\tau$ is a threshold hyperparameter. The reverse cyclic consistency map $M_{1\leftarrow 2}$ is computed with similar procedure.
@@
Figure 3: Visualization of epipolar lines. Our predictions can well follow the ground truth even under large viewpoint changes.
To estimate camera parameters, we use the knowledge taken from previous study [40] and adopt essential matrix module to output rotation $R$ and translation $t$. The essential matrix module is a mapping module that exploits each transformer token from images to a feature that is used to predict $R$ and $t$. This module contains positional encoding, bilinear attention, and dual softmax over attention map $A$. Following the design in Sec. 3.3, we make a modification to replace $A$ with our cost volume, since it acts as a key for emulating 8-point algorithm, such that the better spatial correspondences encoded in $A$ can aid more accurate camera pose estimation. Subsequent to the essential matrix module, we finally regress 6D rotation representations [62] and 3 translation parameters with scales using MLPs.
@@ 3.6 Training Objectives
$\begin{split}\mathcal{L}_{\text{match}}=\sum_{i}M_{1\leftarrow 2}(1-\text{SSIM}(F_{1\leftarrow 2}(I_{2}(i)),I_{1}(i)))\\ +M_{2\leftarrow 1}(1-\text{SSIM}(F_{2\leftarrow 1}(I_{1}(i)),I_{2}(i))).\end{split}$ (1)
Figure 5: Qualitative comparison on ACID.
@@ 4.1 Implementation Details
Our encoder uses ResNet-34, taking 256$\times$256 image as an input, and extracts 3 levels of feature maps with a spatial resolution of 16, 32, and 64. We set $\lambda_{\text{tri}}=0.01$. Our network is implemented using PyTorch [38] and trained with the AdamW [32] optimizer. We set the base learning rate as 2e44-4- 4 and use an effective batch size of 64. The network is trained for 50K iterations, taking approximately 2 days. We exponentially decay the learning rate with $\gamma=0.95$ after every epoch. We train and evaluate all other baselines on the same datasets for fair comparisons. We provide training and evaluation details of ours and our competitors in the supp. material for completeness. The pre-trained weights and codes will be made publicly available at https://ku-cvlab.github.io/CoPoNeRF/.
@@ 4.2 Experimental Settings
We assess the performance of our method on two tasks: novel-view synthesis and relative pose estimation. The latter task serves a dual purpose as it also assesses the quality of our correspondences, a methodological approach that aligns with prior image matching studies. We first compare with established generalized NeRF variants, specifically PixelNeRF [59] and [17], highlighting the complexities of wide-baseline inputs. Subsequently, we compare with existing pose-free generalized NeRF methods, namely DBARF [11] and FlowCAM [45]. For relative pose estimation, the evaluation includes several comparisons: classical matching methods [15, 43, 49] followed by RANSAC [21] and end-to-end pose estimation frameworks [60, 40]. Finally, we conduct a comparative analysis with pose-free generalized NeRF methods [11, 45].
@@
Table 1: Pose estimation performance on RealEstate-10K and ACID. Gray color indicates methods not directly comparable as they supervise correspondence with ground-truth depth; they are included for reference only. We also specify the targeted task for each method. *: FlowCAM [45] results have been updated to rectify an error in the numerical values originally presented. We apologize for any inconvenience caused.
@@
Table 2: Novel view rendering performance on RealEstate-10K and ACID. Gray text indicates methods not directly comparable for their use of ground-truth pose at evaluation. *: FlowCAM [45] results have been updated to rectify an error in the numerical values originally presented.
@@ 4.3 Experimental Results
FlowCAM* [45] + pose loss ($\mathcal{L}_{\text{pose}}$)
@@
Tab. 2 shows quantitative comparisons, whereas Fig. 5 and Fig. 6 show qualitative comparisons. From the results, compared to previous pose-free approaches [45, 11], our approach outperforms them all. Note that we also include results from generalized NeRFs [59, 17] to highlight the complexity and challenge of this task. It’s important to note that while our method may not surpass the state-of-the-art [17], the proximity of our results to it underlines the potential and effectiveness of our approach in the absence of camera pose information.
@@ 4.4 Ablation Study and Analysis
In Tab. 3, we validate the effectiveness of each component within our framework. The baseline in the first row represents a variant equipped with only the feature backbone, renderer, pose head and image reconstruction loss. We then progressively add each component. From the results, we observe clear improvements on performance for every component, demonstrating that they all contribute to the final performance. A particularly illustrative comparison are (I) vs (II) and (II) vs (III), where simply adding each loss leads to apparent improvements, indicating that the process of finding correspondences, learning 3D geometry through rendering and estimating camera all contribute largely to the performance.
@@
Unless our framework achieves more competitive rendering quality, the practicality of our method will be rather limited. In this analysis, we compare our framework with the variants that adopt two separate methods for camera pose estimation and rendering. Specifically, for the first row, we combine pretrained generalized NeRF [17] with an off-the-shelf matching network for pose estiomation. The second row shows the outcomes obtained with [40]. Note that RelPose [60] does not predict translations, and we thus leverage [40]. Table LABEL:tab:fixed reports the results. Summarizing the results, we observe that our approach outperforms the other variants by large margin, highlighting the importance of capturing the underlying synergy between the tasks and the practicality of the proposed approach.
@@ 5 Conclusion
In this work, we have presented a novel unified framework that integrates camera pose estimation, NeRF rendering and correspondence estimation. This approach effectively overcomes the limitations of existing approaches, particularly in scenarios with limited data and complex geometries. Our experimental results, encompassing both indoor and outdoor scenes with only a pair of wide-baseline images, demonstrate the framework’s robustness and adaptability in achieving high-quality novel view synthesis and precise camera pose estimation. Extensive ablation studies further validates our choices and highlight the potential of our method to set a new standard in the this task.

### 08 Apr 2024 v2

# Unifying Correspondence, Pose and NeRF for Generalized Pose-Free Novel View Synthesis

@@ 2 Related Work
3 Unified Framework for Generalized Pose-Free Novel View Synthesis
@@ 3.1 Problem Formulation
Assuming an unposed pair of images $I_{1},I_{2}\in\mathbb{R}^{H\times W\times 3}$ taken from different viewpoints as the input, our goal is to synthesize an image $\hat{I}_{t}$ from a novel view. In this work, we assume camera intrinsics are given, as it is generally available from modern devices [1]. Different from classical generalized NeRF tasks [58, 10, 66], our task is additionally challenged by the absence of camera pose between the input images. To this end, we estimate the relative camera pose between $I_{1},I_{2}$ as ${P}_{2\leftarrow 1}\in\mathbb{R}^{4\times 4},$ consisting of rotation $R\in\mathbb{R}^{3\times 3}$ and translation $T\in\mathbb{R}^{3\times 1}$, and deduce $P_{2\leftarrow t}=P_{2\leftarrow 1}P_{1\leftarrow t}$ with $P_{1\leftarrow t}$ as the desired rendering viewpoint, which are then used in conjunction with the extracted feature maps to compute the pixel color at the novel view by the renderer.
@@ 3.2 Cost Volume Construction
Unlike the previous methods that only consider local receptive fields within their cost volumes [63, 10, 33, 12], we consider all pairs of similarities between features to handle both small and large displacements [14, 15, 28, 29, 30]. Specifically, we compute and store the pairwise cosine similarity between features, obtaining a 4D cost volume $\{C^{l}\}_{l=1}^{L}\in\mathbb{R}^{h^{l}\times w^{l}\times h^{l}\times w^{l}}$, where $L$ is the number of levels.
@@ 3.4 Flow and Relative Pose Estimation
Making use of the cost volume from previous steps, we define a dense flow field, $F(i)$ that warps all pixels $i$ in image $I_{1}$ towards $I_{2}$. We also estimate relative camera pose $P_{2\leftarrow 1}$ from this cost volume as it sufficiently embodies confidence scores and spatial correspondences [9, 46]. To estimate the dense flow map $F_{2\leftarrow 1}$, we can simply apply argmax to find the highest scoring correspondences. While this may be sufficient for image pairs with large overlapping regions, we account for the potential occlusions by computing a confidence map. Specifically, following [40], we obtain a cyclic consistency map $M_{2\leftarrow 1}(i)$ using the reverse field $F_{1\leftarrow 2}$ as an additional input, and check if the following condition is met for consistency: $||F_{2\leftarrow 1}(i)+F_{1\leftarrow 2}(i+F_{2\leftarrow 1}(i))||_{2}<\tau$, where $||\cdot||_{2}$ denotes frobenius norm and $\tau$ is a threshold hyperparameter. The reverse cyclic consistency map $M_{1\leftarrow 2}$ is computed with similar procedure.
@@
Figure 3: Visualization of epipolar lines. We use the relative camera pose to draw epipolar lines based on the points in (a). Our predictions can well follow the ground truth even under large viewpoint changes.
To estimate camera parameters, we use the knowledge taken from previous study [46] and adopt an essential matrix module to output rotation $R$ and translation $t$. The essential matrix module is a mapping module that exploits each transformer token from images to a feature that is used to predict $R$ and $t$. This module contains positional encoding, bilinear attention, and dual softmax over attention map $A$. Following the design in Sec. 3.3, we make a modification to replace $A$ with our cost volume, since it acts as a key for emulating 8-point algorithm, such that the better spatial correspondences encoded in $A$ can aid more accurate camera pose estimation. Subsequent to the essential matrix module, we finally regress 6D rotation representations [69] and 3 translation parameters with scales using MLPs.
@@ 3.6 Training Objectives
$\begin{split}\mathcal{L}_{\text{match}}=\sum_{i}M_{1\leftarrow 2}(i)(1-\text{SSIM}(F_{1\leftarrow 2}(I_{2}(i)),I_{1}(i)))\\ +M_{2\leftarrow 1}(i)(1-\text{SSIM}(F_{2\leftarrow 1}(I_{1}(i)),I_{2}(i))),\end{split}$ (1)
where $\phi(\cdot,\cdot)$ yields the SSIM score between two comparative inputs.
@@ 4.1 Implementation Details
Our encoder uses ResNet-34, taking 256$\times$256 image as an input, and extracts 3 levels of feature maps with a spatial resolution of 16, 32, and 64. We set $\lambda_{\text{tri}}=0.01$. Our network is implemented using PyTorch [44] and trained with the AdamW [38] optimizer. We set the base learning rate as 2e44-4- 4 and use an effective batch size of 64. The network is trained for 50K iterations, taking approximately 2 days. We exponentially decay the learning rate with $\gamma=0.95$ after every epoch. We train and evaluate all other baselines on the same datasets for fair comparisons. We provide training and evaluation details of ours and our competitors in the supp. material for completeness.
@@ 4.2 Experimental Settings
We assess the performance of our method on two tasks: novel-view synthesis and relative pose estimation. The latter task serves a dual purpose as it also assesses the quality of our correspondences, a methodological approach that aligns with prior image matching studies. We first compare with established generalized NeRF variants, specifically PixelNeRF [66] and [20], highlighting the complexities of wide-baseline inputs. Subsequently, we compare with existing pose-free generalized NeRF methods, namely DBARF [12] and FlowCAM [52]. For relative pose estimation, the evaluation includes several comparisons: matching methods [18, 50, 56] followed by 5-point algorithm [43] and RANSAC [24] and end-to-end pose estimation frameworks [67, 46]. Finally, we conduct a comparative analysis with pose-free generalized NeRF methods [12, 52].
@@
Table 1: Pose estimation performance on RealEstate-10K and ACID. Gray color indicates methods not directly comparable as they supervise correspondence with ground-truth depth; they are included for reference only. We also specify the targeted task for each method.
@@
Table 2: Novel view rendering performance on RealEstate-10K and ACID. Gray text indicates methods not directly comparable for their use of ground-truth pose at evaluation.
@@ 4.3 Experimental Results
FlowCAM [52] + pose loss ($\mathcal{L}_{\text{pose}}$)
@@
Tab. 2 shows quantitative comparisons, whereas Fig. 5 show qualitative comparisons. From the results, compared to previous pose-free approaches [52, 12], our approach outperforms them all. Note that we also include results from generalized NeRFs [66, 20] to highlight the complexity and challenge of this task. It’s important to note that while our method may not surpass the state-of-the-art [20], the proximity of our results to it underlines the potential and effectiveness of our approach in the absence of camera pose information.
@@ 4.4 Ablation Study and Analysis
In Tab. 3, we validate the effectiveness of each component within our framework. The baseline in the first row represents a variant equipped with only the feature backbone, renderer, pose head and image reconstruction loss. We then progressively add each component. From the results, we observe clear improvements on performance for every component, demonstrating that they all contribute to the final performance. A particularly illustrative comparison are (I) vs (II) and (II) vs (III), where simply adding each loss leads to apparent improvements, indicating that the process of finding correspondences, learning 3D geometry through rendering and estimating camera all contribute largely to the performance. However, it is also notable that PDC-Net+. + [20] reports lower rotation and translation angular errors. This can be attributed to the use of classical solvers [24, 43] that are known to output more precise transformations given a sufficient numbe of correspondences [47].
@@
Unless our framework achieves more competitive rendering quality, the practicality of our method will be rather limited. In this analysis, we compare our framework with the variants that adopt two separate methods for camera pose estimation and rendering. The results are reported in Table LABEL:tab:fixed. Specifically, for the first row, we combine pretrained generalized NeRF [20] with an off-the-shelf matching network for pose estimation. The second row shows the outcomes obtained with [46]. Note that RelPose [67] does not predict translations, and we thus leverage [46]. Summarizing the results, we observe that our approach outperforms the other variants by large margin, highlighting the importance of capturing the underlying synergy between the tasks and the practicality of the proposed approach.
@@ 5 Conclusion
In this work, we have presented a novel unified framework that integrates camera pose estimation, NeRF rendering and correspondence estimation. This approach effectively overcomes the limitations of existing approaches, particularly in scenarios with limited data and complex geometries. Our experimental results, encompassing both indoor and outdoor scenes with only a pair of wide-baseline images, demonstrate the framework’s robustness and adaptability in achieving high-quality novel view synthesis and precise camera pose estimation. Extensive ablation studies further validates our choices and highlight the potential of our method to set a new standard in this task.
Acknowledgement
This research was supported by the MSIT, Korea (IITP-2023-2020-0-01819, RS-2023-00222280).